# Few-Class Arena: A Benchmark for Efficient Selection of Vision Models and Dataset Difficulty Measurement

### v1

@@ A.7 Extended Few-Class Similarity Benchmark (FC-Sim) Results
Overall, a key observation is that the general trend among all ten datasets unveils the inverse relationship between similarity and the number of classes. Specifically, image similarities, which act as proxy of inverse subset difficulty score increases as the number of classes $N_{CL}$ decreases. This reveals that similarity plays a more important role in the Few-Class Regime than for datasets with more classes. Therefore, for real applications with few classes, simply downscaling a model blindly without considering class similarity may yield a model selection with sub-optimal efficiency. We propose, therefore, that image similarity must be taken into consideration for existing scaling laws [26, 12, 88]. To that end, Few-Class Arenas developed to facilitate future research in this direction.
@@
Effect of ResNet Scales on Similarity. We present the details of FCA-Sim results of the ResNet family in different scales in the Few-Class Regime of ImageNet1K, specifically ($N_{CL}\in\{2,3,4,5,10,100\}$) shown in Fig. 19. In particular, we analyze the relationship between each full and sub-Model’s Top-1 accuracy and SimSS by Pearson correlation coefficient (PCC) denoted as $r$ in the plots. The ResNet family scales from ResNet18 to ResNet152. We experiment both CLIP (red in the first two rows) and DINOv2 (blue in the last two rows) as similarity base functions.
@@
Transformer vs Similarity. We present the details of Vision Transformer (ViT) [75] Base performance and SimSS using CLIP and DINOv2 in the subsets of $N_{CL}\in\{2,3,4,5,10,100\}$ in ImageNet1K, shown in Fig. 20. An overview of the results indicates low correlation between Top-1 accuracy and SimSS ($r\leq 0.20$). Our tool Few-Class Arena helps discover the new challenge of using the vanilla ViT in the Few-Class Regime, as ViTs are originally designed to scale up [88]  when large training datasets in the order of billion samples are available. Few-Class Arena also identifies the importance of scaling down considering similarity.
(a) Full
(b) Full
(c) Sub
(d) Sub
Figure 20: Top-1 Accuracy of ViT Base vs SimSS on ImageNet1K. $N_{CL}\in\{2,3,4,5,10,100\}$.
A.8 Experiments Compute Resources
Experiments are conducted in two internal clusters with the following specifications: (1) 8 NVIDIA RTX A5000 24GB, AMD EPYC 7513 32-Core Processor 882GB RAM and (2) 8 NVIDIA TITAN Xp 12GB, Intel(R) Xeon(R) CPU E5-2650 v4 @ 2.20GHz, 126GB RAM. When GPUs in two clusters are fully utilized, training 10 models in 9 datasets takes 2 weeks; obtaining one experiment result of FC-Full usually takes less than 1 minute since it only involves inference without training; obtaining one experiment result of FC-Sub takes around 2 days on average depending on the size of subset and model, which includes both training and testing; computing the SimSS in the Few-Class Regime for ten datasets takes around three weeks.

### v2

@@ A.10 Extended Few-Class Similarity Benchmark (FC-Sim) Results
Overall, a key observation is that the general trend among all ten datasets unveils the inverse relationship between similarity and the number of classes. Specifically, image similarities, which act as proxy of inverse subset difficulty score increases as the number of classes $N_{CL}$ decreases. This reveals that similarity plays a more important role in the Few-Class Regime than for datasets with more classes. Therefore, for real applications with few classes, simply downscaling a model blindly without considering class similarity may yield a model selection with sub-optimal efficiency. We propose, therefore, that image similarity must be taken into consideration for existing scaling laws (Kaplan et al., 2020; Rae et al., 2021; Zhai et al., 2022). To that end, Few-Class Arena is developed to facilitate future research in this direction.
@@
Effect of ResNet Scales on Similarity. We present the details of FC-Sim results of the ResNet family in different scales in the Few-Class Regime of ImageNet1K, specifically ($N_{CL}\in\{2,3,4,5,10,100\}$) shown in Fig. 19. In particular, we analyze the relationship between each full and sub-model’s Top-1 accuracy and SimSS by Pearson correlation coefficient (PCC) denoted as $r$ in the plots. The ResNet family scales from ResNet18 to ResNet152. We experiment both CLIP (dash line in the 1st and 3rd rows) and DINOv2 (solid line in the 2nd and 4th rows) as similarity base functions.
@@
A.11 Comparing Various Architectures
We perform experiments on CNNs (ResNet18, 50) and Vision Transformer-Base (ViT-B). Results are summarized in Table 7. The overall trend of ViT-B exhibits behavior consistent with the observations for ResNet described in Fig. 1.
MT $N_{CL}$ ResNet18 ResNet50 ViT-B
[…]
F 1000 69.90 0 76.55 0 82.37 0
F 10 66.68 4.372 73.76 3.675 79.52 2.704
F 5 65.68 9.680 72.16 9.302 79.12 8.146
F 2 62.80 14.18 70.60 14.67 78.00 14.47
S 10 91.88 1.640 91.48 2.265 82.16 3.508
S 5 93.68 1.213 92.96 3.170 89.60 2.227
S 2 96.80 1.304 94.80 2.168 95.60 1.949
Table 7: Performance results for different models and configurations on ImageNet1K. $N_{CL}\in\{2,5,10,1000\}$, MT: Model Type, F: Full model, S: Sub-model.
A.12 Extension to Object Detection and Segmentation
To assess the generalization of the few-class properties to Object Detection (OD) and Segmentation (Seg), we conducted validation experiments using YOLOv8. The procedure is consistent with the method outlined in Section 3.3 and 3.4. Specifically, for a specific $N_{CL}$ (2 in this example), we randomly sample the $N_{CL}$ classes from the full dataset of COCO, where each consists of five subsets with seed numbers from 0 to 4. We performed experiments with 2, 5, 80. The YOLOv8-nano model was chosen since we focus on efficiency. Image size of 320x320 was used. Model performance was evaluated using the standard metric, mean average precision at an IoU threshold of 0.5 (mAP@50). Table 8 summarizes our results:
MT $N_{CL}$ OD Seg
[…]
F 80 0.405 0.195 0.378 0.200
F 5 0.456 0.090 0.435 0.098
FT 5 0.488 0.069 0.465 0.082
S 5 0.503 0.069 0.474 0.084
F 2 0.488 0.161 0.475 0.180
FT 2 0.505 0.127 0.457 0.159
S 2 0.538 0.106 0.482 0.152
Table 8: Performance results for different tasks and configurations on COCO. $N_{CL}\in\{2,5,80\}$, MT: Model Type, OD: Object Detection, Seg: Segmentation, F: Full model, S: Sub-model, FT: Fine-tuned model. Best scores are highlighted in bold. The gray bar indicates sub-models as the primary focus of this research.
We also include a comparison with fine-tuned pre-trained models (denoted as F). These models are initialized with weights from the pre-trained YOLOv8-nano on the full dataset with 80 classes, and subsequently fine-tuned for a limited number of epochs (five in this example). We conclude that OD and Seg task exhibit similar observations to those we make in Fig. 1.
A.13 Experiments Compute Resources
Experiments are conducted on two internal clusters with the following hardware specifications: (1) 8 NVIDIA RTX A5000 GPUs (24GB), an AMD EPYC 7513 32-Core Processor, and 882GB of RAM; and (2) 8 NVIDIA TITAN Xp GPUs (12GB), an Intel(R) Xeon(R) CPU E5-2650 v4 @2.20GHz, and 126GB of RAM. When GPUs in two clusters are fully utilized, training ten models in nine datasets takes two weeks; obtaining a single experiment result for FC-Full usually takes less than one minute since it only involves inference without training; getting one FC-Sub experiment result takes approximately two days on average depending on the size of subset and model, which includes both training and testing; computing the SimSS in the Few-Class Regime for ten datasets takes around three weeks. Additional experiments are conducted on a gaming desktop with the following hardware specifications: 2 NVIDIA RTX 3090 Ti GPUs (24GB), an AMD Ryzen 5 3600 6-Core Processor, and 78.5GB of RAM.